OpenStudy (anonymous):

The equation of a circle whose center is at (4, 0) and radius is length 2√(3) is

OpenStudy (anonymous):

Can anyone help me please ?

OpenStudy (anonymous):

Do you have answer chioces to choose from

OpenStudy (anonymous):

@Summi45618 (x - 4)² + y² = 2√3 (x - 4)² + y² = 12 (x + 4)² + y² = 12

OpenStudy (anonymous):

Well it could not be c because of the (x +4 that needs to be (x-4

OpenStudy (anonymous):

Ok so we know what the x and y axis are right.....SO x-axis is 4 and the y-axis is 0 Sp try to fill in the variable and tell me what answer you get

OpenStudy (anonymous):

And which answer choice works for you

OpenStudy (anonymous):

Is it "B"? @Summi45618

OpenStudy (anonymous):

Um...Good Try but not quiet....Here I have an example let me show you

OpenStudy (anonymous):

Use the standard equation for a circle of radius r with its centre at (p, q) Then (x - p)² + (y - q)² = r² Substituting the values given in the question - I've assumed the radius is meant to be 2√3 but if this is wrong the basic process is still the same. (x - 4)² + (y - 0)² = [2√3]² = 12 (x - 4)² + y² = 12 The LHS can be expanded if required. x² - 8x + 16 + y² = 12........and then simplified x² - 8x + y² = -4

OpenStudy (anonymous):

Does this example make sense?

OpenStudy (anonymous):

I'm sorry. I am trying to break it down so I can apply it to the question I have but my program has a time thingie on it so I keep going back inbetween the two sites to make sure it doesnt timeout because it wont let me log back in. But im trying to understand :/ @Summi45618

OpenStudy (anonymous):

OH OH IS IT A? @Summi45618

OpenStudy (anonymous):

Correct!!!!!!!!!

OpenStudy (anonymous):

YAYYYY THANK YOU ! Thats only number two out of 30 . Omg I suck @Summi45618

OpenStudy (anonymous):

@DisCrazyGinger Thank you !!!!!